The resonant LC circuit inside your radio contains a coil with an inductance of L = 0.496 mH. By turning the radio knob you adjust the capacitor to C = 34.4 pF. What is the wavelength of the radio waves your radio is receiving?
2.462×102 m

You are correct.
Your receipt no. is 159-9737 Help: Receipt
Previous Tries

What is the speed of the radio waves received by your radio?

PLEASE HELP!!
Answer #1

we know,
v = f*λ

Now λ = 2.462 * 10^2 m
Freq can be calculated as,
f = 1/(2π*sqrt(LC))
f = 1/(2π*sqrt(0.496*10^-3*34.4*10^-12))
f = 1.218 * 10^6 Hz

v = 1.218 * 10^6 *  2.462 * 10^2
v = 299871600 m/s
v = 2.998 * 10^8 m/s